# MtgrBoost: Boosting Large-scale Generative Recommendation Models in Meituan

### v1

@@ 4.2 Embedding Table Merging
Industrial recommendation systems usually merge multiple separate embedding tables into a larger and unified table. Merging tables offers two key advantages. First, the merged embedding table is stored as a contiguous memory block, which avoids memory fragmentation issues. Second, the merged table significantly enhances lookup efficiency, as multiple table lookup operations are fused into a single operation. However, we observe critical limitations in TorchRec’s table merging implementations. The merging process requires developers to manually configure each embedding table individually, which becomes labor-intensive for massive tables.
@@
Automated Merging Table. To address the above issue, we propose an automated table merging framework. We design a unified feature configuration interface, FeatureConfig, which enables automated merging by defining parameters for each feature (e.g., feature name, embedding dimensions, and lookup tables). MTGRBoost then generates merging strategies, such as combining tables with identical embedding dimensions. To support dynamic embedding table merging, we design HashTableCollection, which generates a collection of hash tables using feature configurations and performs pooling computations as needed. Finally, MTGRBoost retrieves embeddings via sequence IDs, with cross-table lookups managed through predefined lookup tables. Developers need only specify required features and MTGRBoost automatically registers features and executes table merging without manual intervention in implementation details.
@@
Previous Solution. Merging table may lead to numerical overlap in the ID space of different tables. For example, both user ID and item ID might have the same ID value in the whole merged table. To address this issue, recommendation systems typically employ a row-wise offset mechanism. As shown in Figure 7 (a), a unique row offset value is assigned to each embedding table (i.e., offset 100 for table 2), calculated by summing the number of rows from all preceding embedding tables. This offset is then added to the original ID to generate a globally unique ID, which is used to retrieve the corresponding embedding from the merged table.
@@
Our Solution. The dynamic table merging also encounters the issue of ID numerical overlap. However, because the number of rows in dynamic tables cannot be accessible in advance and thus fixed row offsets cannot be directly applied. To address this issue, we allocate the largest possible offset for each dynamic table to prevent ID overflow. We still use row offsets to calculate globally unique IDs, however, these offsets are dynamically calculated based on the number of feature tables. Specifically, we use bitwise operations to combine the original ID with a feature table identifier to create a globally unique ID. We begin by computing the number of bits needed for encoding feature table indices as $k=\lceil\log_{2}(m+1)\rceil$, using the high $k$ bits of the 64-bit integer space as feature identifier bits. We ensure the offset is a positive integer by setting the highest bit to 0, while the remaining $(64-1-k)$ bits are used to calculate the maximum number of row in the hash table. As illustrated in Figure 7 (b), we use 2 identifier bits to distinguish between three tables (since $\lceil\log_{2}(3+1)\rceil=2$), while calculating the maximum row capacity of hash table as $2^{61}$ through utilization of the remaining 61 bits (64-1-2). Consequently, the offset values for “Table 2” and “Table 3” are configured as $2^{59}$ and $2^{60}$, respectively, derived through successive halving of the maximum row capacity. Formally, the calculation formula for the ID $x$ in the $i$-th feature table is as follows:
$\mathrm{ID}=(i\ll(63-k))\mid x,$ (8)
@@
where $\ll$ denotes shift left operation, $\mid$ is bitwise OR.
4.3 Embedding Table Lookup Acceleration
For industrial recommendation systems, embedding tables are typically divided into shards distributed across multiple devices. Exchanging embedding vectors across devices requires two all-to-all collective communications for each table lookup: ID communication and embedding communication. ID communication involves exchanging feature IDs to be queried, while embedding communication entails exchanging the queried embeddings. However, a sequence batch may contain numerous duplicate feature IDs, and limited bandwidth can render ID communication and embedding communication as performance bottlenecks. Additionally, duplicate feature IDs result in redundant reads of the same embedding vector, significantly increasing memory access pressure.
@@ 4.3 Embedding Table Lookup Acceleration
Figure 8: A running example of two-stage ID deduplication. We only visualize the ID deduplication process on GPU 0 and omit the embedding communication.
@@
Two-stage ID Deduplication. To address the issues mentioned, Figure 8 illustrates a two-stage ID deduplication process aimed at reducing communication latency and lookup frequency. In the first stage, latency is decreased in two critical areas: feature ID communication and embedding communication. Devices within nodes deduplicate feature IDs before exchanging them, resulting in reduced ID communication latency. More importantly, deduplication significantly alleviates embedding communication latency. If IDs are redundant, devices will transmit multiple duplicate embeddings, as they must return the corresponding embedding vectors based on the received feature IDs. Despite optimization efforts in the first stage, the all-to-all ID exchange inherently reintroduces duplicate IDs across devices due to distributed communication patterns. The second deduplication stage removes these newly generated duplicates, thereby minimizing the frequency of table lookups. Notably, the primary objective of the two-stage ID uniqueness operation is to reduce embedding communication latency, as ID communication has a relatively minor impact and hash table lookups are inherently efficient.
5 Additional Optimizations
5.1 Sequence Balancing
@@
User sequences naturally exhibit a long-tail distribution: a small subset of highly active users generate exceptionally long sequences, while the majority produce relatively short ones. Previous deep recommendation models (DRMs) are constrained by architectural limitations that enforce sequence truncation and padding for length alignment. In contrast, generative recommendation models (GRMs) preserve complete user sequences, thereby achieving superior accuracy. This advantage arises because truncation risks removing semantically critical tokens (i.e., an item in the sequence), even when separated by many intermediate interactions. For example, a user’s early phone purchase history may influence subsequent accessory recommendations despite temporal distance. Furthermore, truncation alters the fundamental semantics of sequential dependencies.
@@
Figure 9: The visualization of imbalanced computational load. We train the MTGRBoost on 8 GPUs and report the maximum and minimum GPU computation times across steps 0–20. The shaded region indicates GPU idle time.
@@
Initial Design. FlashAttention [3] processes sequences by dividing them into chunks, where each chunk independently computes local attention. The final results are iteratively merged to cover the full sequence. This chunk-based design inherently supports variable-length sequences, motivating our initial approach of directly computing complete user sequences with flash attention. We then randomly sample a fixed number of sequences to form training batches and distribute them across devices. However, this straightforward batching approach resulted in significant efficiency challenges. Figure 9 illustrates the range of GPU computation times per training step. The shaded region in the figure reveals prolonged synchronization delays (up to 25.8 ms) across devices, resulting in resource underutilization and reduced training efficiency. The primary cause is the unequal distribution of sequence length: within a single training step, devices process token quantities differing by as much as 40,000. To mitigate this issue, we propose a novel dynamic sequence batching method to balance computational loads across devices.
Input: Target token count $N$, input chunks $\mathcal{C}$

### v2

@@ 3.2. Embedding Table Merging
Industrial recommendation systems usually merge multiple embedding tables into a unified table, which improves efficiency by fusing multiple lookup operations into one (jiang2019xdl; zhang2022picasso). However, the merging process requires developers to manually configure each embedding table, which becomes labor-intensive for massive tables. To address this issue, we propose an automated table merging technique.
@@
Automated Merging Table. We design a unified feature configuration interface, FeatureConfig, which enables automated table merging by defining parameters for each feature (e.g., feature name, embedding dimensions, and lookup tables). MTGenRec then generates merging strategies, such as combining tables with identical embedding dimensions. To support dynamic embedding table merging, we introduce a HashTableCollection that uses feature configurations and performs pooling computations as needed. Developers only need to specify the required features, and MTGenRec automatically registers them and executes the table merging without requiring manual intervention. The detailed workflow and introduction are provided in Appendix B.3.
@@
Previous Solution. Table merging can accelerate the lookup process, however, it introduces numerical overlap in the ID spaces of different tables. For example, user IDs and item IDs might share the same value in the merged table. TorchRec (ivchenko2022torchrec) addresses this issue using a row-wise offset mechanism for static table merging. In particular, each embedding table is assigned a unique row offset, which is calculated by summing the number of rows from all preceding tables. This offset is then added to the original ID to generate a globally unique ID, which is used to retrieve the corresponding embedding from the merged embedding table.
@@
Our Solution. Dynamic table merging also faces ID overlap issues. However, the row counts of dynamic tables are unpredictable in advance, which precludes the use of fixed offsets. To resolve this, we dynamically calculate offsets based on the number of tables to prevent ID overflow. Specifically, we use bitwise operations to combine the original ID with a feature table identifier to create a globally unique ID. We first compute the number of bits needed for encoding feature table indices as $k=\lceil\log_{2}(m+1)\rceil$, using the high $k$ bits of the 64-bit integer space as feature identifier bits. The highest bit is set to 0 to ensure the offset is a positive integer, while the remaining (641k)(64-1-k) bits are used to determine the maximum number of rows in the hash table. For example, if we use 2 identifier bits to distinguish between 3 tables (since $\lceil\log_{2}(3+1)\rceil=2$), the remaining 61 bits (64-1-2) are used to calculate the maximum row capacity. Consequently, the offset values for “Table 2” and “Table 3” are configured as $2^{59}$ and $2^{60}$, respectively, derived through successive halving of the maximum row capacity. The calculation formula of the ID $x$ in the $i$-th feature table is expressed as follows:
(8) $\mathrm{ID}=(i\ll(63-k))\mid x,$
@@
where $\ll$ denotes shift left operation, $\mid$ is bitwise OR. A detailed running case is provided in Appendix B.4 for better understanding.
@@
Figure 6. A running example of two-stage ID deduplication. We only visualize the ID deduplication process on GPU 0 and omit the embedding communication.
3.3. Embedding Table Lookup Acceleration
Embedding tables are typically divided into shards distributed across devices. Each table lookup involves two all-to-all collective communications: one for ID exchange and another for embedding vector exchange. The communication cost is directly proportional to the number of features. However, a sequence batch may contain numerous duplicate features IDs, and wasted bandwidth can render the lookup communication as performance bottlenecks. To address this issue, we propose a two-stage ID deduplication technique aimed at reducing communication latency.
@@ 3.3. Embedding Table Lookup Acceleration
Two-stage ID Deduplication. As shown in Figure 6, the first stage involves deduplicating feature IDs on each device before exchanging them, thereby reducing ID communication latency. Despite this optimization, the ID exchange process can reintroduce duplicates due to different devices sending the same ID. Thus, the second deduplication stage is necessary to remove these newly generated duplicates. ID deduplication can significantly alleviate embedding communication latency. If feature IDs are redundant, devices will transmit multiple duplicate embeddings, as they must return the corresponding embedding vectors based on the received feature IDs. Notably, the primary objective of the two-stage ID deduplication is to reduce embedding communication latency, as ID communication has a relatively minor impact.
3.4. Sequence Balancing
@@
In addition to optimizing the sparse model of GRM, we observe that existing methods are still insufficient in optimizing the dense model. User sequences exhibit a long-tail distribution: a small subset of highly active users generate exceptionally long sequences, while the most produce short ones. Previous DRMs are constrained by architectural limitations that enforce sequence truncation and padding for length alignment. In contrast, GRMs require preserving complete user sequences to achieve superior accuracy. This advantage arises because truncation risks removing semantically critical tokens, even when separated by many intermediate interactions.
@@
Figure 7. The visualization of imbalanced computational load. We train the MTGenRec on 8 GPUs and report the maximum and minimum GPU computation times across steps 0–20.
@@
Initial Design. FlashAttention (dao2022flashattention) processes sequences by dividing them into chunks, where each chunk independently computes local attention. The final results are iteratively merged to cover the full sequence. This chunk-based design inherently supports variable-length sequences, motivating our initial approach of directly computing complete user sequences. However, this straightforward batching approach results in significant efficiency challenges. Figure 7 illustrates the GPU computation times per training step. The shaded region in the figure reveals prolonged synchronization delays (up to 25.8 ms) across devices, resulting in resource underutilization and reduced training efficiency. The primary cause is the unequal distribution of sequence lengths, which leads to a load imbalance issue among devices.
Figure 8. Running example of dynamic sequence batching. The number indicates the number of tokens in a sample.
Dynamic Sequence Batching. To address the problem, as shown in Figure 8, we introduce a dynamic sequence batching technique. Specifically, each GPU maintains a buffer for storing sequence samples retrieved from Hive table chunks. We define the target token count as the product of average sequence length (i.e., 600) and the batch size. We first compute token counts per sample in the buffer and calculate their cumulative sums. A binary search algorithm then identifies the optimal partition point where the cumulative sum most closely approaches the target token count. We finally obtain a balanced batch by outputting the first $k$ sequences in the buffer. If the total number of tokens in the buffer falls below the target, the remaining samples are merged into the subsequent buffer. This process iterates until all Hive table chunk samples are consumed. A detailed algorithm is provided in Appendix B.5.
In data parallelism, gradients are typically averaged using All-Reduce operations. However, with dynamic sequence batching, directly averaging gradients may introduce biases, as each device computes gradients based on a varying number of samples. To address this issue, we implement All-to-all communication to synchronize batch sizes across devices, followed by performing a weighted average of gradients proportional to their respective batch sizes. This ensures numerical correctness in both loss computation and gradient update processes.
Discussions. Dynamic sequence batching provides two critical advantages: (1) improved device load balancing via token constraints, and (2) optimized hardware utilization via adaptive batch sizes. A fixed batch size strategy requires conservative configurations to prevent out-of-memory risks from clusters of extremely long sequences. In contrast, dynamic batching maximizes device memory utilization by loading batches near device memory limits. While more complex load balancing methods exist, such as sequence packing based on length distributions (krell2021efficient; kundu2024enhancing) or dynamic programming-based parallelism strategies (li2024demystifying), we ultimately exclude these methods from our MTGenRec. This decision stems from the empirical observation that our lightweight dynamic sequence balancing achieves near-optimal workload balancing under current training conditions, rendering additional complexity unnecessary.
Besides the above optimizations for sparse and dense model, we incorporate a suite of additional techniques to implement a comprehensive GRM training system, including checkpoint resumption, mixed-precision training, gradient accumulation, and operator fusion. A detailed introduction is provided in Appendix C.
Figure 9. The GAUC results of CTR and CTCVR on GRM 4G 1D for TorchRec and MTGenRec, respectively.
4. Experimental Evaluations
4.1. Experiment Settings
[…]
Baselines. We compare MTGenRec with TorchRec (ivchenko2022torchrec), the state-of-the-art GRM training framework proposed by Meta. Actually, our initial design is developed based on TorchRec, and we further implement MTGenRec with the proposed optimizations. We conduct experimental comparisons against TorchRec regarding both model accuracy and efficiency. We exclude alternative baselines as TorchRec is the only PyTorch-native training framework specifically optimized for PyTorch-based GRMs. Although other training systems such as TensorFlow are widely used in the industry, their framework currently cannot support GRM training.
Dataset and Implementation. We conduct experiments on 10 days of user logs from Meituan, with daily log data exceeding 200 million sequences. The average sequence length is 600, reaching a maximum length of 3,000. The detailed dataset statistics are provided in Appendix D.1. All models are trained from scratch using the same data. To optimize sparse and dense features, we employ the Adam (kingma2014adam) optimizer.
Evaluation Metrics. We use Group AUC (GAUC) to evaluate the accuracy of the model on CTR and CTCVR tasks and system throughput to evaluate the training efficiency.
Table 1. Comparison of offline and online performances on the Meituan Take-away application. We report the improved accuracy of GRM compared to DRM.
Offline Metric diff Online Metric diff
CTR GAUC CTCVR GAUC PV_CTR UV_CTCVR
GRM 4G +0.0036 +0.0154 +1.04% +0.04%
GRM 110G +0.0153 +0.0288 +1.90% +1.02%
4.2. Comparisons to Baselines
Accuracy and Business Benefits. Figure 9 presents the CTR and CTCVR GAUC results obtained from training a GRM using MTGenRec and TorchRec. The results show that both MTGenRec and TorchRec ensures model correctness and training stability. To further validate the effectiveness of our GRM, we deploy MTGenRec on the Meituan Take-away application, conducting A/B testing with $2\%$ of the traffic. The comparison baseline is the most advanced online DRM (i.e., DLRM (naumov2019deep)), which has been continuously learning for 2 years. As shown in Table  1, although the volume of training data is significantly lower, the offline and online metrics of GRM still greatly exceed those of the baseline.
Table 2. Time (second) decomposition for GRM 4G 1D and GRM 110G 64D in TorchRec and MTGenRec, respectively.
Model Method Lookup Forward Backward
GRM 4G 1D TorchRec 7.0 (s) 14.8 (s) 50.7 (s)
MTGenRec 3.2 (s) 14.0 (s) 32.6 (s)
GRM 110G 64D TorchRec 95.2 (s) 27.8 (s) 177.7 (s)
MTGenRec 14.4 (s) 19.9 (s) 54.7 (s)
Time Decomposition. Table 2 presents execution times for GRM 4G 1D and 50G 64D over 100 training steps, detailing embedding lookup, forward, and backward phases. The results show that MTGenRec achieves faster execution than TorchRec across all phases. The efficiency improvements from sequence balancing increase substantially with model complexity. Larger embedding dimensions induce significant increases in the lookup and backward phases due to communication overhead and sparse parameter update latency.
Figure 10. Ablation studies for GRM 4G 1D and GRM 110G 1D in MTGenRec. Compared with the baseline, MTGenRec achieves a 1.60 $\times$ to 2.44 $\times$ throughput improvement.
Figure 11. Throughput comparison for disabling and enabling sequence balancing on GRM 4G 1D and 110G 1D. We scale the system from 8 GPUs to 64 GPUs and report the gain values of sequence balancing at the top of figure.
4.3. Evaluation of Individual Designs
Ablation Study. Figure 10 presents an ablation study under 4G and 10G complexity settings by incrementally integrating different techniques in MTGenRec, including merging table, two-stage ID deduplication and sequence balancing. The experimental results demonstrate all these techniques in MTGenRec can enhance system efficiency. Furthermore, we observe that performance gains are more pronounced with higher computational complexity. This trend reveals a positive correlation between MTGenRec ’s efficiency improvements and computational complexity.
Table 3. We report the comparison results in batch size and average GPU memory utilization before and after enabling sequence balancing. Numbers in parentheses denote improvements in GPU memory utilization.
Model Batch size GPU memory utilization
GRM 4G 1D 480 → 496 86.3% → 95.7% (+9.4)
GRM 110G 1D 80 → 116 75.3% → 90.3% (+15.0)
Sequence Balancing Evaluation. We evaluate sequence balancing on GRM 4G and 110G models, scaling from 8 to 64 GPUs. To ensure fairness, batch sizes are adjusted to maintain near-full GPU memory utilization under both enabled and disabled conditions. Figure 11 presents the throughput results, and table 3 reports batch sizes and average GPU memory utilization, yielding four key observations. (1) Sequence balancing significantly improves throughput across all settings (e.g., an average improvements of 26.5% for GRM 110G, peaking at 33.5% on 64 GPUs). (2) Throughput gains scale with GPU count due to reduced synchronization delay from straggler devices. As the number of GPUs increases, the likelihood of encountering long sequences that induce computational delays also rises. (3) Throughput improvements correlate positively with computational complexity. Since sequence processing FLOPs scale quadratically with hidden embedding dimensions, higher computational complexity amplifies load imbalance, which makes sequence balancing increasingly critical. (4) Table 3 shows that sequence balancing maximizes GPU memory utilization, whereas fixed-size batching requires conservative sizing to avoid OOM errors.
Figure 12. Throughput comparisons of two-stage ID deduplication for GRM 4G 1D (upper) and GRM 4G 64D (bottom). The improvements of the current strategy compared to the previous one are reported above the bars.
[…]
Table 4. Throughput comparison for MCH and MTGenRec. OOM denotes out of memory and Gain denotes the throughput improvement. “-” indicates that the comparison cannot be made due to OOM problems.
Complexity Dim. Factor MCH MTGenRec Gain
4G 1D 392,731 579,649 47.59%
8D 260,590 438,190 68.2%
64D 80,857 155,832 92.7%
110G 1D 44,719 68,993 54.28%
8D 28,329 55,929 97.4%
64D OOM 38,575 -
[…]
Figure 13. Comparison of scalability by computational complexity and embedding dimension. We use 8 GPUs configuration as the baseline for speedup ratio. The specific speedup ratio values are marked near the broken line.
4.4. Scalability
[…]
5. Related Work
Systems for DRM Training. Deep Recommendation Models (DRMs) combine sparse embeddings and dense neural networks to model each user-item individually. As DRM is widely used, many systems are developed to accelerate it (lian2022persia; jiang2019xdl; xie2020kraken; miao2021het; jiang2018linear; feng2020atbrg; fan2019mobius; zhao2020distributed; guo2021scalefreectr; smelyanskiy2019zion; krishna2020accelerating; wang2022merlin). Specifically, Persia (lian2022persia) decouples sparse embedding computation from dense model processing, training the embeddings asynchronously to improve throughput and the dense models synchronously to ensure accuracy. PICASSO (zhang2022picasso) leverages the patterns of model architecture and data distribution to accelerate execution with packing, interleaving, and caching optimizations. XDL (jiang2019xdl) compresses the pair-wise communication among GPUs into a tree structure to accelerate embedding lookup and update. ScaleFreeCTR (guo2021scalefreectr) leverages host memory to store massive embedding tables and employs GPU-optimized synchronization to update the embeddings. Zion (smelyanskiy2019zion) and RecSpeed (krishna2020accelerating) mitigate the I/O bottlenecks for embedding communication by deploying additional NICs. However, these systems are typically designed based on specific assumptions about the access patterns of sparse embeddings (e.g., skewed towards frequently accessed embeddings), which limits their applicability compared to our dynamic embedding table. Additionally, these systems are implemented using TensorFlow, making them unsuitable for PyTorch-based GRM training.
Systems for Transformer. To train Transformer efficiently, the focuses are attention computation and parallel strategies (ren2021zero; wang2023zero++; dao2022flashattention; lu2017flexflow; jia2018exploring; huang2019gpipe; shoeybi2019megatron; rasley2020deepspeed; li2024demystifying; wu2024bitpipe; rajbhandari2020zero). FlashAttention (dao2022flashattention) reduces the accesses to slow GPU global memory with tiled computation and online recomputation fusion. FlexFlow (lu2017flexflow) and OptCNN (jia2018exploring) automatically search the optimal parallel strategy for model training by distributing model computation and parameters among the GPUs in different ways. GPipe (huang2019gpipe) employs pipeline parallelism by splitting each mini-batches into micro-batches to interleave the computation and communication across the micro-batches. Megatron-LM (shoeybi2019megatron) and DeepSpeed (rasley2020deepspeed) propose hybrid parallel strategies for large-scale model training by partitioning the model and data across the GPUs. Hydraulis (li2024demystifying) adopts dynamic parallelism for each iteration according to the sequence length distribution to achieve workload balance. MTGenRec enjoys existing attention kernel optimizations for efficient computation but simple data parallel suffices for GRMs since their dense models are usually small. Moreover, MTGenRec optimizes the processing of sparse embeddings, which are not present in most Transformer-based models.
6. Conclusion
We propose MTGenRec, a distributed training system designed to enhance scalability and efficiency of industrial generative recommendation models (GRMs). Our sparse model optimizations include dynamic embedding tables for changeable IDs, automated table merging, and two-stage ID deduplication. For dense models, dynamic sequence batching mitigates workload imbalances. The experiment results demonstrate substantially improved training efficiency without compromising accuracy, paving the way for broader adoption of GRMs in recommendation systems.
Appendix A The Differences between GRM and DRM
Our model builds upon GR (zhai2024actions) by introducing a novel prediction head enhanced through MMoE. This design replaces GR’s MLP with MMoE, which employs task-specific gate networks to dynamically balance multiple objectives, improving accuracy in multi-task scenarios. Compared to MLP’s parameter-intensive tuning, MMoE reduces computational overhead by selectively activating only the top-k expert outputs.
Additionally, our MTGenRec diverges from DRM in batch construction. As illustrated in Figure 14, DLRM’s pairwise batches contain redundant user computations (e.g., ”User 1” features are processed twice). In contrast, MTGenRec employs a sequence-wise approach that consolidates multiple user interactions into a single sample, which ensures each user feature is computed once. Notably, user features contain orders of magnitude more tokens than item features (e.g., 10,000 vs. 100), enabling MTGenRec to significantly accelerate training and improve scalability.
Appendix B Additionally Details for MTGenRec
B.1. Pipeline.
To enhance training efficiency, we maximize parallelism through pipeline technology using three streams: copy, dispatch, and compute. Specifically, the copy stream loads sequences from CPU to GPU, the dispatch stream performs table lookups based on IDs, and the compute stream handles forward computation and backward updates. For instance, while the compute stream executes forward and backward passes for batch $\mathbf{T}$, the copy stream concurrently loads batch $\mathbf{T}+1$ to mask I/O latency. Upon completing backward updates for batch $\mathbf{T}$, the dispatch stream immediately initiates table lookups and communication for batch $\mathbf{T}+1$.
Figure 14. Batch structure of DRM and GRM. “U” denotes users, “P” denotes products. DRM and GRM use a pairwise and sequence-wise batch construction method, respectively.
B.2. Proof
Lemma B.1.
Proof.
$S$ contains no factor of 2 (since $S$ is odd), while $M$ has prime factorization $2^{n}$. Thus, $S$ and $M$ share no common prime factors, proving $\mathrm{gcd}(S,M)=1$. ∎
Proof.
Assume there exist $t_{1},t_{2}\!\in\![0,M\!-\!1]$, $t_{1}\!\neq\!t_{2}$ such that:
(9) $h_{t_{1}}\equiv h_{t_{2}},$
This implies:
(10) $(t_{1}-t_{2})\cdot S\equiv 0.$
(11) $t_{1}-t_{2}\equiv 0.$
Proof.
(12) $\{h_{t}\}_{t=0}^{M-1}=\{0,1,2,...,M-1\}.$
Figure 15. We achieve automatic table merging by two embedding configuration interfaces.
B.3. Interface for Automated Table Merging
In industrial recommendation systems, it is common practice to consolidate multiple embedding tables into a larger unified table. Frameworks like TorchRec’s EmbeddingCollection, for instance, support merging tables with identical embedding dimensions. However, we observe significant limitations in existing GRM approaches regarding table merging. The aforementioned table merging process requires developers to manually configure each embedding table individually, which becomes excessively laborious when dealing with massive numbers of embedding tables.
To address this challenge, we design an automated table merging interface. Specifically, Figure 15 illustrates our unified feature configuration interface, FeatureConfig, which enables automatic table merging by defining a configuration entry for each feature based on its name, embedding vocabulary size, embedding dimension, and lookup table. Subsequently, MTGenRec reads data from the database according to feature names and generates a table merging strategy (e.g., merging tables with identical embedding dimensions). To support the merging of dynamic embedding tables, we design HashtableCollection, which generates a collection of hash tables based on the embedding configurations (e.g., feature name, embedding dimension) within the FeatureConfig list and performs pooling operations if required. Finally, MTGenRec retrieves the corresponding embeddings for sequence IDs from the embedding layer. Since the embeddings for certain features might be scattered across multiple tables, we implement multi-table lookup for a single feature via lookup tables. Throughout this workflow, developers only need to configure the required features, with all other details including feature registration and table merging implementation handled automatically by MTGenRec.
B.4. Running Example for Dynamic Offset
As shown in Figure 16 (a), a unique row offset value is assigned to each embedding table (i.e., offset 100 for table 2), calculated by summing the number of rows from all preceding embedding tables. In contrast, as illustrated in Figure 16 (b), we use 2 identifier bits to distinguish between three tables, while calculating the maximum row capacity of hash table as $2^{61}$ through utilization of the remaining 61 bits.
Input: Target token count NN, input chunks $\mathcal{C}$